Circular No. 6405

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ab IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 18.5) located at

α = 15^{h}21^{m}09^{s}.0, δ = +27^{o}55′23′′ (equinox 2000.0), which is 10′′ west and 15′′ north of the center of a galaxy evidently identified as Abell 2065-356 by Postman et al. (1988, A.J. 95, 267), but which apparently is not the host galaxy (see report by Garnavich et al., below), a conclusion supported by N. Reid’s finding that there is extended emission underlying SN 1996ab that is apparently not connected with the larger system. SN 1996ab was found on a IIIa-J plate taken by K. M. Rykoski and J. D. Mendenhall on May 12 UT with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. No object appears at the position of SN 1996ab on original Sky Survey prints; Reid further reports that there is no object at this position on Sky Survey plates exposed in 1986 May and June.

P. Garnavich, A. Riess, P. Challis, and R. Kirshner, Harvard-Smith-sonian Center for Astrophysics, write: “Spectra obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 22.4 UT show that this is a type-Ia supernova near maximum light. The Si II 635.5-nm absorption is observed at 692 nm, which implies a redshift of z = 0.13 for the supernova (assuming an expansion velocity of 10 000 km/s). SN 1996ab is probably not associated with the Abell Cluster 2065 (z = 0.06), but rather with a fainter galaxy background to the cluster.”

CYGNUS X-1 S. N. Zhang, W. S. Paciesas, B. A. Harmon, G. J. Fishman, and

D. Crary, Marshall Space Flight Center, NASA, report for the Compton Gamma Ray Observatory BATSE team: “Cygnus X-1 has been decreasing rapidly in intensity in the band 20-200 keV, as observed by BATSE, roughly coincident in time with the increased 2- to 12-keV flux reported by ASM/RXTE (see IAUC 6404). The flux changed from 1.0 ± 0.02 Crab units (20–100 keV) on May 11 to 0.5 ± 0.02 on May 19–20. The energy spectrum also becomes softer during this time (the Optically Thin Thermal Bremsstrahlung model temperature being from 80 to ∼ 50 keV). The r.m.s. fraction (0.01–0.488 Hz) also decreased from a typical value of 8 ± 0.5 to 5 ± 0.5 percent during this period.”

1996 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green